# RoboMatch: A Unified Mobile-Manipulation Teleoperation Platform with Auto-Matching Network Architecture for Long-Horizon Tasks

### v2

@@ Abstract
This paper presents RoboMatch, a novel unified teleoperation platform for mobile manipulation with an auto-matching network architecture, designed to tackle long-horizon tasks in dynamic environments. Our system enhances teleoperation performance, data collection efficiency, task accuracy, and operational stability. The core of RoboMatch is a cockpit-style control interface that enables synchronous operation of the mobile base and dual arms, significantly improving control precision and data collection. Moreover, we introduce the Proprioceptive-Visual Enhanced Diffusion Policy (PVE-DP), which leverages Discrete Wavelet Transform (DWT) for multi-scale visual feature extraction and integrates high-precision IMUs at the end-effector to enrich proprioceptive feedback, substantially boosting fine manipulation performance. Furthermore, we propose an Auto-Matching Network (AMN) architecture that decomposes long-horizon tasks into logical sequences and dynamically assigns lightweight pre-trained models for distributed inference. Experimental results demonstrate that our approach improves data collection efficiency by over 20%, increases task success rates by 20–30% with PVE-DP, and enhances long-horizon inference performance by approximately 40% with AMN, offering a robust solution for complex manipulation tasks.
@@ I INTRODUCTION
Imitation learning from human demonstrations offers an efficient way for robots to acquire skills, allowing humans to teach various tasks. Mobile teleoperation platforms are ideal for complex whole-body coordination tasks like opening doors or moving objects. Current systems mainly follow two approaches: decoupled platforms [1, 2] require two operators for mobility and manipulation separately. While simplifying individual control, dual-user operation may cause efficiency loss and synchronization issues. Integrated platforms face either usability limitations (e.g., Mobile ALOHA’s push-pull control [3]) or flexibility constraints (e.g., HOMIE’s fixed cockpit [4]). These limitations hinder high-quality demonstration collection and restrict imitation learning’s effectiveness in complex tasks. Thus, developing convenient and flexible teleoperation platforms has become crucial for advancing robot imitation learning.
@@
Imitation learning has gained significant attention in robotics due to its practical applicability. Strategies like ACT [5] and Diffusion Policy [6] enable autonomous decision-making through end-to-end learning from demonstrations and environmental perception. Such perception includes visual [7] and proprioceptive sensing [8]. While visual representations encode scene information, learning robust visual features for robotics remains challenging. Attention mechanisms like the Efficient Multi-scale Attention Module (EMA) [9] enhance visual perception but focus only on spatial features, ignoring frequency-domain information. Proprioceptive accuracy critically affects motion control precision. During delicate tasks like assembly or grasping, insufficient data and error accumulation often cause end-effector pose deviations, leading to failure. Improving proprioception and reducing error propagation are thus key challenges in embodied intelligence.
@@
As robotic intelligence advances, task planning, as a critical bridge connecting high-level instructions with low-level execution, faces growing challenges. Traditional rule-based methods often fall short in long-horizon, multi-step tasks within open environments. Emerging large language models (LLMs) [10] offer a promising alternative by parsing abstract instructions into executable task sequences through natural language understanding and logical reasoning. When integrated with visual-language models (VLMs) [11, 12], robots achieve end-to-end intelligence from perception to planning. For example, encoding 3D scenes into structured representations enables VLMs to generate physically constrained strategies [13]. Breakthrough end-to-end vision-language-action (VLA) [14, 15, 16] architectures further unify perception, decision-making, and execution in a single model. However, critical challenges remain: constraint satisfaction struggles with complex multi-step tasks; VLAs exhibit limited long-horizon reasoning and hallucinations [17]; while substantial network architectures cause latency and error accumulation, hindering real-world deployment [18]. Overcoming these bottlenecks is essential to enhance the reliability and adaptability of robotic task planning.
@@ II-A Teleoperation for Mobile Manipulation
Robot Teleoperation [19, 20] enables real-time remote control via intermediary systems and is widely used in industrial, medical, and collaborative scenarios. However, mobile manipulation research remains nascent, especially in achieving efficient mobile base-arm coordination [10]. While early systems adopted master-slave architectures [21], recent advances like Mobile ALOHA [3] and HOMIE [4] integrate mobility and manipulation. Some approaches leverage high-end hardware (e.g., motion capture [22], exoskeletons [4]) or adapt static interfaces (e.g., VR [23], keyboard [24]) to mobile platforms. Yet, As shown in Table I, most solutions still lack fine multi-DoF control, task generalization, operational integration, and long-term deployment capabilities, limiting their applicability in complex mobile manipulation tasks requiring high precision and integration.
@@
Teleop System
Modality
Action
Space
Cockpit
Bimanual
Exo-
Teleop
Whole-Body
Teleop
Wild
@@ II-B Imitation Learning
Imitation learning has become a key approach in robotics [28, 29], enabling robots to acquire manipulation strategies for complex tasks from expert demonstrations [30, 31]. Among these methods, behavior cloning (BC) serves as a foundational technique that directly maps observations to actions [32], offering strong transferability and proving particularly suitable for policy learning based on teleoperation data. Recent advances integrate imitation learning with deep learning frameworks, adopting network strategies such as Transformer [5, 33] and Diffusion [6] for autonomous decision-making in complex continuous tasks. Notably, diffusion policies have demonstrated remarkable generalization in handling intricate robotic tasks and are widely adopted in VLA large models [16, 34, 35].
Despite their strong task adaptability, diffusion models still exhibit noticeable limitations in motion generation precision, especially in fine manipulation tasks (e.g., folding clothes) that rely solely on joint angles and simple spatial image observations. Cumulative errors during inference, coupled with insufficient visual perception and proprioception, significantly impede precise decision-making and efficient robot control [36].
@@ II-C Long-Horizon Inference in Robotics
Recent studies show that LLM-based planning agents offer substantial advantages in long-horizon task decomposition and execution [10, 37]. The rapid development of large language models has made natural language instruction prompting a major trend in robot task planning. Recently, VLMs have demonstrated broad potential in this domain, with research mainly following two technical pathways: some approaches employ VLMs as task planners or formulate them as constraint satisfaction problems [11]. For instance, the ReKep system [13] uses VLMs to map 3D environmental keypoints into numerical cost functions and optimizes grasping poses by minimizing collision risk and instability metrics. Other approaches explore integrated VLA models that combine VLMs with action heads, providing a feasible framework for end-to-end robot learning, as exemplified by OpenVLA [14] and $\pi$0.5 [15].
@@ III-A RoboMatch Hardware Design and Manipulation
In terms of design, we adopted an inverted master-slave manipulator layout to form an exoskeleton-based teleoperation configuration, enhancing operational intuitiveness by aligning with the operator’s limb motion scale. Dedicated foot pedals with embedded IMUs were developed for feet-controlled base movement. Each pedal measures rotation around the Pitch axis, mapping the angle to velocity commands: the left pedal controls forward/backward motion with speed proportional to tilt, while the right pedal enables precise left/right rotation using the same mapping logic.
@@ III-B The Model Framework of PVE-DP
$P_{\text{fused}}=\text{Concat}(X_{\text{qpos}},X_{\text{quat}}),$ (3)
At the algorithmic level, we introduce a Frequency Enhanced EMA (FE-EMA) module and deeply integrate it with the visual backbone of the diffusion policy. This module enables efficient fusion of spatial and frequency domain features, thereby constructing spatio-frequency visual feature representations.
@@
Given an input feature map $X$, its wavelet transform is expressed as Eq. (4), achieved by constructing a set of scaled and translated wavelet basis functions $\psi_{j,k}(x)$ and scaling functions $\varphi_{j,k}(x)$
@@
For a 2D feature map $X$ the wavelet transform can be implemented using 2D filters. It decomposes the feature map along rows and columns, yielding the following coefficients: $\text{DWT}(X)$ = $(cA,cH,cV,cD)$,where cAcA denotes the low-frequency approximation coefficients, capturing global structural information;$cH$ represents the horizontal detail components, extracting horizontal edge features; $cV$ corresponds to the vertical detail components, capturing vertical edge features; and $cD$ indicates the diagonal detail components, characterizing diagonal edge features. The Haar wavelet filter is employed to extract both the low-frequency component $cA$ and the high-frequency components $cH$, cVcV and $cD$.
@@
In Eq. (5), $X_{\text{space}}$ represents the spatial domain feature generated through an attention mechanism and group normalization process. It primarily captures local regional information extracted from the input features. This feature is combined with the frequency-domain feature $X_{\text{freq}}$ to ensure that the model captures more comprehensive multi-scale information. Such an approach enables dynamic balancing between spatial and frequency domain features, effectively enhancing the model’s performance.
@@
During the denoising process, PVE-DP continuously refines the predicted action sequence by incorporating additional wrist perception data from robotic arms IMUs, along with spatio-frequency features extracted by the FE-EMA network, enabling precise motion execution in fine manipulation tasks.
@@ III-C Auto-Matching Network (AMN) Architecture
During inference, the operator provides a command via voice or text. The robot observes the scene and sends the image and instruction to the VLM. The AMN architecture decomposes the task into sequential steps, and then assigns each step a pre-trained specialized policy network for execution. The PVE-DP policy is matched to perform robotic manipulation. After each subtask, the system automatically selects the next network until the entire task is completed.
@@ IV-B Ablation Experiments of PVE-DP
To thoroughly evaluate the performance of PVE-DP, we compare it against state-of-the-art baseline models in the field. The experiments include four tasks: two simulated tasks from the ALOHA benchmark (cube transfer and bimanual insertion), and two real-robot tasks (wiping table and sorting workpiece). IMU modules are integrated into the dual-arm robotic system in simulation to support algorithm validation.
@@ IV-C RoboMatch Performance Experiment
To evaluate RoboMatch systematically, this study uses a split-type data collection platform for baseline comparison through a dual-metric (efficiency and quality) analysis. The test set includes 100 demonstrations—containing Real tasks from Table II and custom tasks—with duration varying from 4s to 32s at 4s intervals. We measure the one-time collection success rate for both platforms and train a unified policy on the obtained data to evaluate quality differences via inference testing.
@@ V CONCLUSIONS
In this paper, we propose RoboMatch, a collaborative, precise, portable, and flexible platform designed for unified mobile-manipulation teleoperation of the whole robot, which enables accurate execution of operational tasks and supports long-horizon inference in real-world environments. The proposed RoboMatch platform significantly reduces the complexity and human cost of teleoperation. The PVE-DP method enhances fine manipulation capabilities. The AMN framework improves operational flexibility, enabling seamless integration of different skills while increasing the success rate in long-horizon tasks. In the future, we will focus on optimizing the RoboMatch cockpit to enhance operator comfort and convenience during data collection, refining the PVE-DP algorithm to further improve robotic manipulation performance, and reducing gaps between subtasks in the AMN framework to enable more intelligent inference.

### v3

@@ Abstract
This paper presents RoboMatch, a novel unified teleoperation platform for mobile manipulation with an auto-matching network architecture, designed to tackle long-horizon tasks in dynamic environments. Our system enhances teleoperation performance, data collection efficiency, task accuracy, and operational stability. The core of RoboMatch is a cockpit-style control interface that enables synchronous operation of the mobile base and dual arms, significantly improving control precision and data collection. Moreover, we introduce the Proprioceptive-Visual Enhanced Diffusion Policy (PVE-DP), which leverages Discrete Wavelet Transform (DWT) for multi-scale visual feature extraction and integrates high-precision IMUs at the end-effector to enrich proprioceptive feedback, substantially boosting fine manipulation performance. Furthermore, we propose an Auto-Matching Network (AMN) architecture that decomposes long-horizon tasks into logical sequences and dynamically assigns lightweight pre-trained models for distributed inference. Experimental results demonstrate that our approach improves data collection efficiency by over 20%, increases task success rates by 20–30% with PVE-DP, and enhances long-horizon inference performance by approximately 40% with AMN, offering a robust solution for complex manipulation tasks. Project website: https://robomatch.github.io
@@ I INTRODUCTION
Imitation learning from human demonstrations offers an efficient way for robots to acquire skills, allowing humans to teach various tasks. Mobile teleoperation platforms are ideal for complex whole-body coordination tasks like opening doors or moving objects. Current systems mainly follow two approaches: decoupled platforms [1] require distinct operations for mobility and manipulation. While simplifying individual control, this divided approach may cause efficiency loss and synchronization issues. Integrated platforms face either usability limitations (e.g., Mobile ALOHA’s push-pull control [2]) or flexibility constraints (e.g., HOMIE’s fixed cockpit [3]). These limitations hinder high-quality demonstration collection and restrict imitation learning’s effectiveness in complex tasks.
@@
Imitation learning has gained significant attention in robotics due to its practical applicability. Strategies like ACT [4] and Diffusion Policy [5] enable autonomous decision-making through end-to-end learning from demonstrations and environmental perception. Such perception includes visual [6] and proprioceptive sensing [7]. While visual representations encode scene information, learning robust visual features for robotics remains challenging. Attention mechanisms like the Efficient Multi-scale Attention Module (EMA) [8] enhance visual perception but focus only on spatial features, ignoring frequency-domain information. Proprioceptive accuracy critically affects motion control precision. During delicate tasks like assembly or grasping, insufficient data and error accumulation often cause end-effector pose deviations, leading to failure.
@@
As robotic intelligence advances, task planning, as a critical bridge connecting high-level instructions with low-level execution, faces growing challenges. Traditional rule-based methods often fall short in long-horizon, multi-step tasks within open environments. Emerging large language models (LLMs) [9] offer a promising alternative by parsing abstract instructions into executable task sequences through natural language understanding and logical reasoning. When integrated with visual-language models (VLMs) [10, 11], robots can transcend simple perception to achieve complex task reasoning. For instance, ReKep [12] leverages VLMs to generate physically grounded cost functions over structured 3D representations, enabling precise, constrained strategy execution. Breakthrough end-to-end vision-language-action (VLA) [13, 14, 15] architectures further unify perception, decision-making, and execution in a single model. However, critical challenges remain: constraint satisfaction struggles with complex multi-step tasks; VLAs exhibit limited long-horizon reasoning and hallucinations [16]; while substantial network architectures cause latency and error accumulation, hindering real-world deployment [17].
@@ II-A Teleoperation for Mobile Manipulation
Robot Teleoperation [18, 19] enables real-time remote control via intermediary systems and is widely used in industrial, medical, and collaborative scenarios. However, mobile manipulation research remains nascent, especially in achieving efficient mobile base-arm coordination. While early systems adopted master-slave architectures [20], recent advances like Mobile ALOHA [2] and HOMIE [3] integrate mobility and manipulation. Some approaches leverage high-end hardware (e.g., motion capture [21], exoskeletons [3]) or adapt static interfaces (e.g., VR [22], keyboard [23]) to mobile platforms. Yet, as shown in Table I, most solutions still lack fine multi-DoF control, task generalization, operational integration, and long-term deployment capabilities, limiting their applicability in complex mobile manipulation tasks requiring high precision and integration.
@@
Teleop System Modality Action Space Cockpit Bimanual Exo- Teleop Whole-Body Teleop Wild
@@ II-B Imitation Learning
Imitation learning has become a key approach in robotics [28, 29], enabling robots to acquire manipulation strategies for complex tasks from expert demonstrations [30, 31]. Among these methods, behavior cloning (BC) serves as a foundational technique that directly maps observations to actions [32], offering strong transferability and proving particularly suitable for policy learning based on teleoperation data. Recent advances integrate imitation learning with deep learning frameworks, adopting network strategies such as Transformer [4, 33] and Diffusion [5] for autonomous decision-making in complex continuous tasks. Notably, diffusion models have demonstrated remarkable generalization in handling intricate robotic tasks and are widely adopted in VLA large models [15, 34, 35].
Despite their strong task adaptability, diffusion models still exhibit noticeable limitations in motion generation precision. Cumulative errors during inference, coupled with insufficient visual perception and proprioception, significantly impede precise decision-making and efficient robot control [36].
@@ II-C Long-Horizon Inference in Robotics
The decomposition and execution of long-horizon tasks have long been a core focus in robotics [37], with recent studies demonstrating that LLM-based planning agents offer substantial advantages in this domain [9]. The rapid development of large language models has made natural language instruction prompting a major trend in robot task planning. Recently, VLMs have demonstrated broad potential, with research mainly dividing into two technical pathways. In the first pathway, some approaches employ VLMs as task planners or formulate them as constraint satisfaction problems [10]. For instance, the ReKep system [12] uses VLMs to generate numerical cost functions defined over 3D environmental keypoints, optimizing relational constraints across the robot-object system to achieve multi-stage task execution. The second pathway explores integrated VLA models that combine VLMs with action heads, providing a feasible framework for end-to-end robot learning, as exemplified by OpenVLA [13] and $\pi$0.5 [14].
@@ III-A RoboMatch Hardware Design and Manipulation
In terms of design, we adopted an inverted master-slave manipulator layout to form an exoskeleton-based teleoperation configuration, enhancing operational intuitiveness by aligning with the operator’s natural motion mapping. Dedicated foot pedals with embedded IMUs were developed for feet-controlled base movement. Each pedal measures rotation around the Pitch axis, mapping the angle to velocity commands: the left pedal controls forward/backward motion with speed proportional to tilt, while the right pedal enables precise left/right rotation using the same mapping logic.
@@ III-B The Model Framework of PVE-DP
$P_{\text{fused}}=\text{Concat}(X_{\text{qpos}},X_{\text{quat}}).$ (3)
Although forward kinematics can compute end-effector orientation, errors in joint positions inevitably propagate to the estimated quaternion. The IMU is not intended to merely replace FK computation; instead, $X_{\text{quat}}$ provides direct physical orientation feedback, complementing $X_{\text{qpos}}$ at the proprioceptive level.
At the algorithmic level, we introduce a FE-EMA module and deeply integrate it with the visual backbone of the diffusion policy. This module enables efficient fusion of spatial and frequency domain features, thereby constructing spatio-frequency visual feature representations.
@@
Given an input feature map $X$, its wavelet transform follows Eq. (4) via a set of scaled and translated wavelet basis functions $\psi_{j,k}(x)$ and scaling functions $\varphi_{j,k}(x)$:
@@
For a 2D feature map $X$ the wavelet transform can be implemented using 2D filters. It decomposes the feature map along rows and columns, yielding the following coefficients: $\text{DWT}(X)$ = $(cA,cH,cV,cD)$, where $cA$ denotes the low-frequency approximation coefficients, capturing global structural information;$cH$ represents the horizontal detail components, extracting horizontal edge features; $cV$ corresponds to the vertical detail components, capturing vertical edge features; and $cD$ indicates the diagonal detail components, characterizing diagonal edge features.
@@
In Eq. (5), $X_{\text{space}}$ represents the spatial domain feature generated through an attention mechanism and group normalization process. It primarily captures local regional information extracted from the input features. This feature is combined with the frequency-domain feature $X_{\text{freq}}$ to ensure that the model captures more comprehensive multi-scale information. Such an approach enables dynamic balancing between spatial and frequency domain features.
@@
During denoising, PVE-DP continuously refines the predicted action sequence by incorporating additional wrist perception data from robotic arms IMUs, along with spatio-frequency features extracted by the FE-EMA network, enabling precise motion execution in fine manipulation tasks.
@@ III-C Auto-Matching Network (AMN) Architecture
During inference, the operator provides a command via voice or text. The robot observes the scene and sends the image and instruction to the VLM. The AMN architecture decomposes the task into sequential steps grounded in the pre-trained policy library, and then assigns each step a pre-trained specialized policy network for execution. After each subtask, the system automatically transitions to the next network until the entire task is completed.
@@ IV-B Ablation Experiments of PVE-DP
To thoroughly evaluate the performance of PVE-DP, we compare it against state-of-the-art baseline models in the field. The experiments include four tasks: two simulated tasks from the ALOHA benchmark (cube transfer and bimanual insertion), and two real-robot tasks (wiping table and sorting workpiece). To facilitate algorithm validation, IMU modules were integrated into both the simulation platform and the physical RoboMatch test platform.
@@ IV-C RoboMatch Performance Experiment
To evaluate RoboMatch systematically, this study uses a split-type (locomotion and manipulation) data collection platform for baseline comparison through a dual-metric (efficiency and quality) analysis. The test set includes 100 demonstrations—containing real tasks from Table II and custom tasks—with duration varying from 4s to 32s at 4s intervals. We measure the one-time collection success rate for both platforms and train a unified policy on the obtained data to evaluate quality differences via inference testing.
@@ V CONCLUSIONS
In this paper, we propose RoboMatch, a collaborative, precise, portable, and flexible platform designed for unified whole-body mobile-manipulation teleoperation, which enables the accurate execution of operational tasks and supports long-horizon inference in real-world environments. The proposed platform significantly reduces the complexity and human effort of teleoperation. Additionally, the PVE-DP method enhances fine manipulation capabilities, while the AMN framework improves operational flexibility, enabling the seamless integration of diverse skills and increasing success rates in long-horizon tasks. Future work will focus on optimizing the RoboMatch cockpit to enhance operator ergonomics during data collection, refining the PVE-DP method to further elevate robotic manipulation performance, and smoothing the transitions between subtasks within the AMN framework to facilitate more cohesive inference.